### v2

# Simulation as Supervision: Mechanistic Pretraining for Scientific Discovery

Simulation as Supervision: Mechanistic Pretraining for Scientific Discovery
Carson Dudley, Reiden Magdaleno, Christopher Harding, Marisa Eisenberg
(November 11, 2025)
1 Abstract
Scientific modeling faces a tradeoff: mechanistic models provide scientific grounding but struggle with real-world complexity, while machine learning models achieve strong predictive performance but require large labeled datasets and are not interpretable. We introduce Simulation-Grounded Neural Networks (SGNNs), which use mechanistic simulations as training data for neural networks. SGNNs are pretrained on synthetic corpora spanning diverse model structures, parameter regimes, stochasticity, and observational artifacts. Simulation-grounded learning has been applied in multiple domains (e.g., surrogate models in physics, forecasting in epidemiology). We provide a unified framework for simulation-grounded learning and evaluated SGNNs across scientific disciplines and modeling tasks. We found that SGNNs were successful across domains: for prediction tasks, they nearly tripled COVID-19 forecasting skill versus CDC baselines, reduced chemical yield prediction error by one-third, and maintained accuracy in ecological forecasting where task-specific models failed. For inference tasks, SGNNs also accurately classified the source of information spread in simulated social networks and enabled supervised learning for unobservable targets, such as estimating COVID-19 transmissibility more accurately than traditional methods even in early outbreaks. Finally, SGNNs enable back-to-simulation attribution, a form of mechanistic interpretability. Back-to-simulation attribution matches real-world observations to the training simulations the model considers most similar, identifying which mechanistic processes the model believes best explain the observed data. By providing a unified framework for simulation-grounded learning, we establish when and how mechanistic simulations can serve as effective training data for robust, interpretable scientific inference.
2 Introduction
Scientific modeling is often divided between two approaches: theory-driven models, which encode domain knowledge into mathematical equations, and data-driven models, which learn flexible patterns from large datasets. Each has fundamental limitations. Mechanistic models are interpretable and grounded in scientific theory, but often break down in practice due to noise, unobserved variables, or incorrect structural assumptions. Neural networks can capture complex relationships but require large labeled datasets, lack scientific grounding, and often learn spurious patterns that fail when applied to new data. We lack models that are both scientifically grounded and capable of robust generalization from imperfect data.
Recent work has sought to bridge this divide by incorporating mechanistic knowledge into machine learning. These approaches can be understood through the three core components of machine learning problems: the data, the model architecture, and the optimization procedure. Physics-Informed Neural Networks (PINNs) incorporate governing equations into the loss function, constraining optimization to satisfy known physical laws [1, 2]. Neural ODEs embed continuous-time dynamics directly into the architecture, parameterizing derivative functions that are solved during forward passes [3]. Universal Differential Equations (UDEs) combine both approaches, using neural networks to learn unknown terms in otherwise known equations while training on trajectory data [4]. These methods perform well when equations are known and all relevant variables are observable, but can introduce bias when mechanistic models are incomplete or when systems undergo regime shifts.
Another approach is to incorporate mechanistic knowledge into the training data itself: using simulations from mechanistic models to generate synthetic datasets for neural network training. This strategy has emerged across disciplines, driven by different motivations. In physics, simulation-trained networks can serve as fast surrogates for expensive simulators, enabling real-time inference where full simulations are computationally prohibitive [5]. In epidemiology and social science, simulations address data scarcity and enable learning of unobservable quantities [6]. Despite strong empirical results in specific applications, these efforts lack a unified framework for understanding when and why simulation-grounded learning succeeds.
We introduce Simulation-Grounded Neural Networks (SGNNs) as a general framework for training neural networks on mechanistic simulations. While simulation-grounded learning has been applied in isolated contexts, we provide a unified treatment and evaluation across domains. SGNNs are pretrained on synthetic corpora spanning diverse model structures, parameter regimes, stochasticity, and realistic observational artifacts. This approach treats mechanistic models as generators of labeled training data, allowing neural networks to internalize scientific structure without requiring it to be incorporated into architectures or loss functions. This gives a flexible framework compatible with any architecture or optimization method.
SGNNs are aligned with recent work demonstrating that large neural networks can acquire general-purpose inference strategies through synthetic supervision [7, 8, 9]. These models show that abstract generative processes can produce strong reasoning priors. SGNNs build on this foundation but differ fundamentally in being grounded in domain-specific scientific simulations. This yields models that are not only performant, but mechanistically interpretable and aligned with scientific theory. In this sense, SGNNs act as scientific inference engines: trained on simulations, and capable of reasoning under real-world uncertainty.
This approach directly addresses a core limitation of machine learning in science: the inability to infer unobservable quantities. Scientific parameters like the basic reproduction number in an infectious disease outbreak or ecological carrying capacity can’t be measured directly. Traditional methods estimate them by fitting mechanistic models to observed trajectories, a process sensitive to noise, delays, and structural mismatch. Simulation-based inference (SBI) addresses this by using simulation-grounded training to learn posterior distributions over parameters [10, 11, 12], providing a principled approach to Bayesian parameter inference. However, full posterior inference is not always necessary: simulation-grounded learning can also enable simpler approaches such as direct quantile regression on parameters, which can achieve accurate inference with reduced computational complexity.
SGNNs also enable back-to-simulation attribution, a form of mechanistic interpretability. After training, the model retrieves training simulations that it considers most similar to real-world inputs, revealing which mechanistic regimes the model associates with observed data. Because retrieved simulations have known ground-truth parameters, users can directly inspect the model’s implicit beliefs about underlying processes. Unlike post hoc feature attribution methods such as SHAP [13] or LIME [14], back-to-simulation attribution provides process-level explanations: not just which input features mattered, but which scientific mechanisms the model believes are active.
We demonstrate SGNNs across tasks in epidemiology, ecology, chemistry, and social science, showing strong generalization in forecasting, regression, classification, and parameter inference. SGNNs nearly triple the median forecasting skill of the CDC Forecast Hub on early COVID-19 data, despite never seeing real-world disease data. Additionally, SGNNs trained only on human-to-human disease simulations outperform specialized models on dengue (a vector-borne disease) by 24%, demonstrating the robustness of the approach. In chemical reaction yield prediction, SGNNs reduce residual variance by one-third, and in ecological forecasting, SGNNs retain high accuracy in high-dimensional settings where task-specific neural nets fail completely.
These results establish SGNNs as a general framework for scientific machine learning, demonstrating when and how mechanistic simulations can serve as effective training data for robust, interpretable inference across disciplines. [^1] Code and data available at https://github.com/carsondudley1/SGNNs.
3 The Simulation-Grounded Neural Network (SGNN) Framework
Figure 1: Simulation-Grounded Neural Networks (SGNNs) use mechanistic simulations as synthetic supervision for scientific modeling. Top: SGNNs are trained on synthetic datasets generated by mechanistic simulators spanning diverse model structures and parameter regimes. Real-world observational artifacts—such as noise, bias, and missingness—are explicitly simulated to mimic real-world data. Neural networks are trained on this synthetic corpus to learn latent mechanistic patterns, enabling generalization to real-world data without retraining. Bottom: SGNNs unify the strengths of mechanistic and data-driven models. They encode scientific theory during training, generalize out-of-the-box without tuning, enable supervised learning for unobservable scientific quantities, support cross-task modeling across domains, and provide process-level interpretability through back-to-simulation attribution.
3.1 Conceptual Framework
SGNNs train neural networks on synthetic data generated by mechanistic simulators. Let $S(\theta)$ denote a stochastic simulator parameterized by $\theta$, which generates trajectories or time series $x_{i}=S(\theta_{i})$. To capture a wide range of system behaviors, we sample $\theta_{i}\sim P(\theta)$ from a task-specific distribution that spans multiple model classes, parameter regimes, and observational settings.
Each simulation yields one or more training examples $x_{i}$ paired with targets $y_{i}=T(\theta_{i})$, where $T(\cdot)$ defines the labeling function (e.g., future trajectories, final states, or latent scientific quantities). Importantly, simulations are passed through an observation model that adds realistic artifacts such as noise, censoring, delays, or reporting bias. The result is a synthetic dataset $\mathcal{D}_{\text{syn}}=\{(x_{i},y_{i})\}$ that reflects both mechanistic diversity and realistic surveillance scenarios.
A neural network $f_{\phi}$ is then trained via supervised learning to predict $y_{i}$ from $x_{i}$. Because each input is paired with exact labels, including latent scientific quantities, and the training set spans a wide range of mechanistic regimes and observational conditions, the model learns to internalize structure that generalizes beyond any single model or parameter setting. SGNNs are architecture-agnostic: any neural network with sufficient capacity may serve as $f_{\phi}$. In practice, we use domain-specific backbones such as temporal transformers, graph networks, or hybrid models (Appendix B).
This simulation-grounded approach supports direct supervision for both observable and unobservable targets, and enables out-of-the-box generalization to real-world data.
3.2 Key Capabilities
Simulation-Grounded Neural Networks (SGNNs) address core limitations of both mechanistic and data-driven modeling approaches. The framework offers five key capabilities:
  • Mechanistic grounding. SGNNs learn from diverse, theory-driven simulations that encode plausible system behaviors, enabling the SGNN to internalize scientific structure.
  • Robustness to misspecification. By training on an ensemble of models spanning multiple structures, parameter regimes, and noise processes, SGNNs develop flexible representations that generalize across regimes—even when real-world dynamics deviate from any one simulator.
  • Supervised learning for unobservable targets. SGNNs enable direct supervision on latent scientific quantities—such as transmission rates or ecological carrying capacities—by training on synthetic datasets where ground truth is known but inaccessible in real-world data.
  • Cross-task generalization. The SGNN approach can be applied across forecasting, regression, classification, and inference tasks.
  • Mechanistic interpretability. Through back-to-simulation attribution, SGNNs identify which simulated regimes they internally associate with real-world inputs, offering insight into the underlying dynamics the model believes are active, and enabling process-level explanations.
4 Applications and Results
4.1 Overview of Evaluation Strategy
We evaluate Simulation-Grounded Neural Networks (SGNNs) across forecasting, inference, regression, and classification tasks spanning multiple scientific domains. Our experiments are designed to assess four core capabilities: (i) generalization to real-world prediction under limited or noisy data, (ii) supervised learning for unobservable scientific parameters via synthetic supervision, (iii) applicability across domains and task types, and (iv) the role of mechanistic grounding in driving model performance.
4.2 SGNNs Enable Robust Infectious Disease Forecasting Across Regimes
4.2.1 SGNNs Outperform CDC COVID-19 Forecasting Models
We first evaluated SGNNs on real-world COVID-19 mortality forecasting during the early pandemic, using weekly death counts from all 50 U.S. states, Washington D.C., and Puerto Rico [15]. This setting poses major challenges: limited early data, rapidly shifting transmission dynamics, evolving interventions, and substantial reporting noise.
The SGNN was pretrained entirely on large-scale mechanistic simulations, which incorporated a wide range of epidemic dynamics, including stochastic super-spreading, policy interventions, seasonality, underreporting, and surveillance delays (Appendix A). No real-world or COVID-specific data were used in pretraining.
We assessed performance using forecasting skill, defined as the percentage improvement in mean absolute error (MAE) relative to a naive baseline that predicts the most recent value:
$\text{Skill}=100\times\left(1-\frac{\text{MAE}_{\text{model}}}{\text{MAE}_{\text{baseline}}}\right)$ (1)
where the baseline prediction is $\hat{y}_{t+h}=y_{t}$ for all forecast horizons $h$. This is a stringent metric in infectious disease forecasting. The mean skill across all models in the CDC Forecast Hub was negative (–1.0) [16]. For comparison, we benchmarked the SGNN against both the median skill (13.0) and best model’s skill (34.0) of Forecast Hub models.
The SGNN outperformed the baselines (Figure 2). Compared to Chronos [17], a state-of-the-art foundation model pretrained on generic time series, the SGNN achieved a roughly 12-fold improvement in skill, underscoring the importance of mechanistic grounding. The SGNN also nearly tripled the Forecast Hub’s median skill and slightly exceeded its best model performance. Finally, the SGNN outperformed a physics-informed neural network (PINN) baseline based on SEIR equations (skill = –112), which performed poorly in our tests in the presence of noise and data sparsity, consistent with prior observations [2]. This demonstrates that incorporating domain knowledge into the data rather than constraining the output of a neural network to match specific equations may be a better strategy in the presence of large uncertainty and minimal real-world data.
Figure 2: SGNNs outperform state-of-the-art baselines in real-world disease forecasting and reveal the importance of mechanistic grounding. (A) SGNNs achieve 35.3% forecasting skill on early COVID-19 mortality, almost tripling the CDC Forecast Hub median and exceeding its best model—despite using no real COVID-19 data. (B) SGNNs produce accurate forecasts with calibrated uncertainty across diverse real-world locations, including Michigan, New York, Texas, and California. (C) Across all U.S. states, SGNNs consistently achieve lower error (lighter) than Chronos and PINNs, showing robust generalization from synthetic pretraining. (D) On dengue, a domain with fundamentally different transmission dynamics, SGNNs outperform both PINNs and statistical models in a fully zero-shot setting—demonstrating robustness to mechanistic misspecification. (E) Replacing mechanistic simulations with neural simulators causes performance collapse, highlighting the necessity of mechanistic fidelity. (F) Restricting pretraining to simple SEIR models leads to overconfident exponential forecasts, confirming that mechanistic diversity and surveillance realism are essential for robust downstream performance.
4.2.2 SGNNs Forecast Accurately Despite Mechanistic Misspecification
Mechanistic models—whether traditional compartmental models or hybrid approaches like PINNs—rely on structural assumptions that may not hold in practice. Real-world systems often deviate from these assumptions, leading to brittle forecasts when the model form is even slightly wrong. This challenge is especially acute when key mechanisms (e.g., transmission routes or feedback loops) are misrepresented or omitted. To assess SGNNs under such conditions, we tested their forecasting performance with different mechanisms from those seen during pretraining.
We evaluated forecasting performance on real-world dengue outbreaks in Brazil [18]. Dengue transmission involves mosquito vectors and mechanisms that are distinct from the human-to-human transmission used in SGNN pretraining simulations. Importantly, SGNNs were evaluated in a fully zero-shot setting: the model had no exposure to dengue data or vector-borne transmission processes during training.
To benchmark performance, we compared SGNNs against two baselines. The first was a PINN-style model trained directly on the dengue data using standard SEIR equations, imposing a rigid, incorrect mechanistic prior [2]. The second was a purely statistical model based on exponential smoothing (ETS), which had access to the dengue data but incorporated no mechanistic assumptions.
SGNNs reduced forecasting error by 57% relative to the PINN, demonstrating robustness under structural misspecification. SGNNs also outperformed the ETS model by 24%, despite operating entirely zero-shot.
These findings underscore a strength of SGNNs: by training on a broad ensemble of mechanistic simulations spanning multiple model structures and parameter regimes, SGNNs learn flexible, scientifically grounded representations that can generalize when structural elements differ from those seen in training. This result suggests that SGNNs do not merely memorize surface patterns, but instead internalize scientific principles that extend across mechanistic boundaries. SGNNs may be learning abstract representations that capture core features of disease spread—such as delays, feedback, and depletion—that hold across different mechanistic forms.
4.3 Realistic, Diverse Mechanistic Grounding is Essential for Generalization
To assess the importance of mechanistic grounding in SGNN training, we performed two ablation studies that isolate the contributions of both model realism and mechanistic diversity (Figure 2).
First, we replaced our simulation pipeline with a neural network trained directly on COVID-19 case data. This neural model acted as a simulator: it generated synthetic outbreaks by recursively forecasting forward from real historical contexts. We then trained a model with the same architecture as our SGNN directly on this simulated data. Despite the model having access to high amounts of simulated data, it was significantly outperformed by the SGNN trained on mechanistic simulations, yielding over twice the error (Figure 2e). This experiment demonstrates that increasing data alone is insufficient. Mechanistic grounding is critical for stable and generalizable learning.
Second, we restricted the simulation corpus to classical SEIR models, removing all other mechanisms (e.g., asymptomatic spread, seasonality, super-spreading, stochasticity) and observational noise (e.g., underreporting, delays). These simulations assumed fixed parameters and homogeneous mixing, and omitted interventions or surveillance artifacts—factors critical for capturing real-world disease dynamics. This ablation isolates the impact of simplification: the underlying mechanistic structure remains broadly correct, but key features of realistic outbreaks are missing. As a result, SGNNs pretrained solely on SEIR data failed to learn these dynamics and produced systematically biased forecasts. For instance, during early COVID-19, they projected unmitigated exponential growth in locations where lockdowns had substantially curbed transmission (Figure 2f). Performance dropped significantly relative to full SGNN pretraining, underscoring that both mechanistic diversity and observational realism are necessary for robust downstream generalization.
4.4 SGNNs Generalize Across Domains and Tasks
4.4.1 SGNNs Forecast Multispecies Ecological Dynamics Where Neural Nets Fail
We evaluated SGNNs on ecological forecasting using two real-world datasets: (i) annual lynx and hare population data from the Hudson’s Bay Company (1821–1934), a low-dimensional predator-prey system [19], and (ii) multivariate butterfly species abundances from the UK Butterfly Monitoring Scheme (UKBMS), which tracks annual fluctuations for dozens of species [20], posing a high-dimensional forecasting challenge.
Our SGNNs were pretrained on ecological simulations representing biological mechanisms and observational noise. For lynx-hare dynamics, training data were generated using stochastic Rosenzweig-MacArthur predator-prey models with demographic noise [21]. For butterfly forecasting, simulations included logistic growth, weak interspecific competition, shared environmental noise, and observation processes such as negative binomial sampling. See Appendix A for full simulator details.
SGNNs achieved strong forecasting skill on both lynx and hare populations. As shown in Figure 3A (left), SGNNs outperformed statistical and mechanistic baselines, with forecasting skill of 48.0 for lynx and 40.2 for hare, higher than the 20.1-32.4 range for all baselines. These results highlight SGNNs ability to generalize to cyclic ecological systems from limited data.
On higher-dimensional forecasting, SGNNs maintained robust forecasting skill as the number of butterfly species increased from 4 to 16 (Figure 3A, right), outperforming task-specific neural networks, which failed entirely in high dimensions. By 16 species, SGNN retained positive forecasting skill above 19.0, while baseline neural nets fell below zero. SGNNs can enable high-dimensional forecasting where standard approaches break down.
Figure 3: SGNNs generalize across scientific domains and task types. (A) Ecological forecasting: SGNNs outperform task-specific neural networks on both low-dimensional predator-prey systems (hare and lynx) and high-dimensional multispecies forecasting from the UK Butterfly Monitoring Scheme. SGNNs maintain forecasting skill as the number of species increases, while baselines degrade sharply. (B) Chemical yield prediction: SGNNs reduce residual variance by one-third compared to state-of-the-art models trained directly on the Suzuki-Miyaura reaction dataset. Right: predicted vs. actual yield for SGNN (blue) and baseline (orange) models, showing tighter clustering around the identity line for SGNN. (C) Diffusion source identification: SGNNs accurately infer the source of information spread in partially observed cascades. Left: top-$k$ accuracy comparison shows SGNNs outperform the Rumor Center method. Right: predicted source probabilities for a representative cascade, with node color and size indicating model-assigned likelihood. SGNN correctly assigns high probability to the true source despite missing observations.
4.4.2 SGNNs Boost Chemical Yield Prediction with Minimal Pretraining
We evaluated SGNNs on the nanomole-scale Suzuki-Miyaura reaction [22] dataset developed by Perera et al. [23], comprising 5,760 reactions spanning diverse combinations of ligands, bases, solvents, and aryl halides under automated flow chemistry. The task is challenging due to high-order interactions, failure-prone edge cases, and wide heterogeneity in reaction yields.
State-of-the-art data-driven models trained directly on this dataset leave 15% of the variance unexplained on held-out test data [24]. SGNNs, pretrained for only a few minutes on high-fidelity chemical reaction network simulations generated via a hybrid simulation pipeline (Appendix A), reduced residual variance to 10%, a one-third reduction. This improvement reflects SGNN’s ability to internalize mechanistic dependencies such as catalyst-substrate incompatibilities, reaction failures, and heteroscedastic noise that purely empirical models fail to capture from limited data alone. A major weakness of purely data-driven models was their inability to distinguish failed reactions, which often led to overconfident but inaccurate predictions in low-yield regimes (Figure 3B, right). By explicitly encoding failure probabilities into our synthetic simulator, SGNNs learned to recognize reactions likely to fail, leading to sharper, more calibrated predictions across the full yield distribution. As shown in Figure 3B, SGNNs yield more accurate predictions with lower residual variance and better alignment with ground-truth yields.
4.4.3 SGNNs Accurately Infer the Source of Diffusion Cascades
We evaluated SGNNs ability to identify the origin of diffusion in networks, a central challenge in computational social science, epidemiology, and security analytics. Given a noisy, partially observed cascade, the model must infer which node initiated the spread.
SGNNs were trained on synthetic cascades generated from the Independent Cascade model on Barabási–Albert networks [25], with 20% of infected nodes randomly masked to simulate incomplete observation. Despite this noise, SGNNs achieved high accuracy: 82.6% top-1 accuracy and over 84.9% top-20 accuracy, versus 58.8% and 72.1% for the Rumor Center method [26]. As shown in Figure 3C, SGNNs substantially outperform baselines on top-$k$ accuracy and produce spatially coherent, well-calibrated predictions across the network, even under missing data.
The SGNN does not approach 100% top-1 accuracy because the true source is masked 20% of the time, meaning that in one out of every five examples, the correct answer is effectively hidden. Yet the SGNN still achieves over 80% accuracy, implying that it sometimes identifies the true source even when that node was never observed to be impacted. While some advanced inference methods can handle unobserved sources under specific assumptions [27], SGNNs achieve this performance in a general setting without explicit source modeling, demonstrating an advance over heuristic or equation-based approaches.
4.5 SGNNs Enable Accurate Inference of Unobservable Parameters
Estimating the basic reproduction number ($R_{0}$)—the expected number of secondary infections generated by a single infectious individual in a fully susceptible population—is central to epidemic response and modeling [28, 29, 30]. However, $R_{0}$ is inherently unobservable, and traditional estimation methods rely on simplified assumptions often violated in practice, especially under data quality issues such as underreporting, delays, and early outbreak noise, as seen in early COVID-19.
To evaluate SGNNs, we pretrained on a diverse set of simulated outbreaks across multiple epidemiological models, with ground-truth $R_{0}$ values computed analytically [31]. These simulations incorporated realistic noise and observation artifacts to mimic surveillance conditions (Appendix A).
We first tested the SGNN on synthetic outbreaks, comparing it to two baselines: (i) SIR-based maximum likelihood estimation (MLE), and (ii) exponential growth-based methods [32]. SGNN achieved an MSE of 1.83 and MPE of 18.6%, outperforming MLE (10.4, 44.5%) and growth-based estimates (8.1, 35.7%) (Figure 4), demonstrating robustness to noise and model misspecification.
We next applied the SGNN to early real-world COVID-19 data (Feb 20-Mar 21, 2020) from U.S. states and cities. During this period, empirical estimates of $R_{0}$ were highly uncertain and often underestimated due to underreporting, inconsistent testing, and data delays. Many early methods assumed exponential growth and homogeneous mixing, neglecting asymptomatic spread and reporting lags.
The SGNN, trained solely on synthetic data, produced $R_{0}$ estimates from early COVID-19 data that were more consistent with later estimates once reporting had improved. For example, early estimates for New York City placed $R_{0}$ from 1.0-2.5 [33], yet subsequent studies using more complete data revised the estimate upward to over 5 [34, 35]. The SGNN, using only data from the early February-March period, correctly produced a high $R_{0}$ estimate for NYC (6.14), consistent with later, more complete analyses, versus 2.35 for exponential growth fitting. A full table of $R_{0}$ estimates is provided in Appendix E.
These results demonstrate that simulation-grounded training enables supervised learning for scientific quantities that are unobservable in real data and highly sensitive to noise and modeling assumptions. SGNN’s robustness to data imperfections and its ability to implicitly infer complex transmission dynamics from short, noisy time series make it well-suited for real-time epidemic inference.
Figure 4: SGNNs accurately infer unobservable parameters and provide mechanistic interpretability via back-to-simulation attribution. (A) SGNNs infer a high reproduction number ($R_{0}=6.14$) for New York City from early COVID-19 case data (Feb–Mar 2020), aligning with estimates from more complete datasets. Traditional methods underestimated early transmission due to underreporting and simplifying assumptions. (B) SGNNs achieve significantly lower mean squared error (left) and mean percentage error (right) in $R_{0}$ estimation compared to maximum likelihood estimation (MLE) and exponential growth-based methods. (C) Back-to-simulation attribution retrieves the 50 most similar synthetic outbreaks to a real-world input and visualizes the distribution of their underlying mechanistic parameters. For Michigan COVID-19 mortality, the retrieved simulations align closely with true estimates for hospitalization rate, population size, asymptomatic transmission rate, and hospitalization fatality rate (dashed lines), confirming that the SGNN’s internal representations encode meaningful mechanistic structure.
4.6 Back-to-Simulation Attribution Enables Mechanistic Interpretability
Beyond predictive performance, SGNNs enable back-to-simulation attribution, allowing real-world inputs to be mechanistically interpreted by identifying which simulated regimes the model deems most similar. Let $f_{\phi}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ denote the trained SGNN encoder that maps input time series to a $d$-dimensional embedding space. For a real-world observation $x$, we retrieve simulations $\{x_{i}\}$ from the training set $\mathcal{D}_{\text{syn}}$ by computing cosine similarity between $f_{\phi}(x)$ and $f_{\phi}(x_{i})$, selecting the top-$k$ most similar embeddings.
This approach differs from direct time series comparison in two ways. First, retrieval in the fixed $d$-dimensional embedding space avoids the curse of dimensionality that would arise from comparing raw high-dimensional time series or parameter spaces. The embedding dimension remains constant regardless of forecast horizon or model complexity. Second, the retrieved simulations reflect the model’s learned similarity rather than time series pattern matching. The encoder $f_{\phi}$ has been trained to map trajectories to representations that are predictive of outcomes. Thus, simulations with similar embeddings are likely the ones the model treats as mechanistically similar for the prediction task.
To perform this attribution, we passed the Michigan COVID-19 mortality time series through the trained SGNN encoder and extracted its embedding. We then computed cosine similarity between this embedding and all simulation embeddings in the synthetic training set. The top-50 most similar simulations were then analyzed to examine the distribution of their underlying mechanistic parameters.
Figure 4 shows violin plots of four key parameters across the top-50 retrieved simulations: case-hospitalization rate, hospitalization fatality rate, asymptomatic probability, and population size. These retrieved values align closely with known or plausible values for COVID-19 in Michigan, capturing the approximate population size (median population 10.1m, versus 10.14 actual), hospitalization and death probabilities consistent with empirical estimates, and realistic asymptomatic proportions. This alignment confirms that the SGNN grounds its predictions in simulations that resemble real-world system dynamics.
Back-to-simulation attribution enables a different form of interpretability than feature attribution and parameter inference methods. Unlike SHAP or LIME [13, 14], which identify which input features influenced a prediction, attribution reveals which mechanistic processes the model believes generated the data. Unlike simulation-based inference approaches that aim to provide statistically rigorous parameter estimation, back-to-simulation attribution explains the model’s predictions by analyzing which mechanistic scenarios it considers similar to the observed data. This provides a domain-aligned window into the model’s predictive reasoning while maintaining computational efficiency.
5 Discussion
SGNNs provide a unified framework for incorporating mechanistic knowledge into machine learning through training data. By generating synthetic datasets from mechanistic simulators, this approach enables neural networks to learn both system dynamics and the observational processes that shape real-world data, including noise, underreporting, and delays.
Simulation-grounded learning bridges mechanistic and empirical modeling by embedding scientific structure through examples rather than architectural constraints or loss functions. This approach is compatible with any architecture and optimization method, supporting transfer learning, fine-tuning, and straightforward integration with existing deep learning infrastructure. Because simulations span diverse model structures and parameter regimes, SGNNs can be robust to structural mismatch between training simulations and real-world systems.
Training on simulated data provides ground-truth labels for quantities that are unobservable in real settings, like transmission rates or carrying capacities that lack direct measurements. This enables approaches like simulation-based inference (SBI), which use supervised learning for parameter inference. Our results show that various approaches are possible within this framework: from full Bayesian posterior estimation, established with SBI, to simpler methods like quantile regression, which can achieve accurate parameter inference with reduced computational complexity.
Back-to-simulation attribution links real-world inputs to training simulations the model considers similar, revealing which mechanistic regimes the model associates with observed patterns. Because each retrieved simulation has known ground-truth parameters and generative processes, users can inspect what the model believes is happening mechanistically. For example, if a model retrieves simulations with high asymptomatic transmission when shown an outbreak with few reported cases, this suggests the model interprets the pattern as indicating substantial hidden spread. Unlike feature attribution methods that identify which input variables influenced a prediction, back-to-simulation attribution explains which scientific processes the model believes are active. This provides actionable insight. If retrieved simulations consistently show specific parameter regimes or structural features, practitioners can validate these mechanistic hypotheses against domain knowledge or use them to guide intervention strategies.
Our results suggest that simulation-grounded learning can capture generalizable representations of scientific structure. Models trained only on human-to-human transmission accurately predicted vector-borne dengue dynamics, indicating that neural networks can internalize abstract principles that transfer across mechanistically distinct settings. This points toward the potential for simulation-pretrained foundation models in science, though substantial work remains to understand the scope and limits of such generalization.
Despite these strengths, SGNNs have important limitations. First, performance depends fundamentally on simulation quality: if key mechanisms are absent from the training simulators, the model cannot learn them. While training on diverse simulator ensembles provides robustness to individual model misspecification, systematic gaps across all training simulations will propagate to predictions. Second, generating comprehensive simulation corpora can be computationally expensive, particularly for high-fidelity mechanistic models. Though this is a one-time cost and inference remains fast, it may be prohibitive for extremely expensive simulators. Third, the sim-to-real gap remains a challenge. Even with realistic noise models and observational artifacts, systematic differences between synthetic and real data can hurt performance. Also, we lack theoretical guarantees about when SGNNs will generalize. While our results show promising zero-shot transfer (e.g., human-to-human models predicting vector-borne disease), we cannot predict a priori which cross-domain generalizations will succeed. Finally, back-to-simulation attribution reveals what the model believes is happening mechanistically, but that can be wrong. Retrieved simulations reflect learned associations that, while often informative, can be incorrect. Understanding when simulation-grounded learning succeeds and when it fails remains an important direction for future work.
SGNNs demonstrate that mechanistic simulations can serve as effective training data across diverse scientific domains, providing a framework for when and how this approach succeeds. By positioning simulation-grounded learning within the broader landscape of hybrid modeling approaches, we establish design principles for combining scientific theory with deep learning flexibility.
Appendix A Simulation Frameworks
A.1 Infectious Disease Simulation Framework
Mechanistic simulator.
To generate synthetic training data for epidemiological forecasting and inference tasks, we developed a universal infectious disease simulator that synthesizes diverse outbreak trajectories across a wide range of plausible regimes. This simulator flexibly instantiates classical compartmental models—such as Susceptible-Infectious-Recovered (SIR), Susceptible-Exposed-Infectious-Recovered (SEIR), and Susceptible-Exposed-Asymptomatic-Infectious-Recovered (SEAIR)—augmented with stochasticity, seasonal variation, demographic turnover, and realistic reporting artifacts. The result is a rich corpus of simulated epidemics that capture both mechanistic diversity and observational realism, serving as the foundation for SGNN pretraining.
Each simulation models a well-mixed population of size $N\sim\text{LogUniform}(5\times 10^{4},5\times 10^{7})$. The compartmental structure is chosen randomly: 70% of simulations include an exposed compartment ($E$), and 50% include an asymptomatic compartment ($A$), resulting in SIR, SEIR, or SEAIR dynamics. The full SEAIR model is governed by:
$\frac{dS}{dt}$ $=-\lambda(t)S+\omega R+\mu N-\mu S,$
$\frac{dE}{dt}$ $=\lambda(t)S-\sigma E-\mu E,$
$\frac{dA}{dt}$ $=p_{A}\sigma E-\gamma A-\mu A,$
$\frac{dI}{dt}$ $=(1-p_{A})\sigma E-\gamma I-\mu I,$
$\frac{dR}{dt}$ $=\gamma I+\gamma A-\omega R-\mu R,$
where parameters are sampled per simulation as follows:
Parameter Distribution Description
$\beta$ $\sim\text{Uniform}(0.10,1.00)$ Base transmission rate per contact
$\gamma$ $\sim\text{Uniform}(0.10,0.33)$ Recovery rate
$\sigma$ $\sim\text{Uniform}(0.20,0.40)$ Latency rate (if exposed state present)
$\omega$ $\sim\text{Uniform}(0.001,0.0075)$ Waning immunity rate (if enabled)
$\mu$ $\sim\text{Uniform}(0,1/365)$ Birth/death rate (if enabled)
$p_{A}$ $\sim\text{Uniform}(0.1,0.7)$ Fraction asymptomatic (if enabled)
$\alpha$ $\sim\text{Uniform}(0.3,1.0)$ Relative transmissibility of $A$ vs. II
Transmission includes super-spreading and seasonal forcing. The time-varying force of infection is:
$\lambda(t)=\beta(t)\cdot\left(\frac{I(t)+\alpha A(t)}{N(t)}\right)\cdot s(t)\cdot\eta(t),$
where $s(t)$ is a sum of 1–4 sinusoidal harmonics (with random amplitudes $\sim\text{Uniform}(0.05,0.20)$ and phases), and $\eta(t)$ is gamma-distributed individual-level overdispersion to model super-spreading. Specifically, $\eta\sim\text{Gamma}(k,k)$ for $k\in[0.1,1.0]$ in a subset of simulations.
To simulate multi-wave epidemics, $\beta(t)$ may shift at 1–5 change points, each with different values sampled from the same $\beta$ distribution. Clinical severity (hospitalization and death probabilities) also varies across waves.
Stochasticity is implemented via discrete-time simulations with daily time steps. For each day, new transitions between compartments are sampled using binomial draws based on the current state and transition probabilities derived from the rates above. This approximates tau-leaping dynamics.
Interventions.
Non-pharmaceutical interventions (NPIs) are included in 25% of simulations. NPIs are triggered when new case counts exceed a sampled threshold $\sim\text{Uniform}(0.001N,0.01N)$. Interventions reduce $\beta(t)$ by a sampled factor $\sim\text{Uniform}(0.2,0.6)$, lasting for $\sim\text{Uniform}(14,120)$ days. They are relaxed when case counts fall below a separate threshold for several days. This simulates lockdowns or behavior change in response to outbreak dynamics.
Demographic turnover and importation.
In 80% of simulations, birth and death processes are active, with Poisson-distributed inflow into $S$ and deaths from all compartments. Importations into $E$ or $I$ occur randomly with daily rate $\sim\text{LogUniform}(0.0001,0.01)$, supporting both isolated and repeatedly seeded outbreaks. If waning immunity is enabled, recovered individuals return to $S$ at rate $\omega$, supporting endemic cycling.
Clinical outcomes.
Hospitalizations and deaths are sampled from symptomatic infections using per-wave probabilities:
  • Hospitalization: $\sim\text{Uniform}(0.02,0.15)$
  • Death given hospitalization: $\sim\text{Uniform}(0.05,0.30)$
Delays from infection to hospitalization and to death are sampled from gamma distributions with: - Hospitalization delay: mean $\sim\text{Uniform}(5,12)$ days - Death delay: mean $\sim\text{Uniform}(14,21)$ days
Observation model.
Realistic surveillance effects are applied to the ground-truth time series. These include:
  • Reporting delay: geometric delay kernels with mode $\sim\text{Uniform}(0,3)$ days
  • Underreporting: case detection probability increases logistically from $\sim\text{Uniform}(5\%,40\%)$ to $\sim\text{Uniform}(25\%,85\%)$ over time, with logistic midpoints drawn uniformly over the simulation window
  • Weekday effects: multiplicative effects $\sim\text{Normal}(1.0,0.05)$ applied to each weekday
  • Noise: lognormal multiplicative noise applied to cases (high), hospitalizations (medium), and deaths (low)
Outputs.
Each simulation returns:
  • True and reported daily time series for infections, hospitalizations, and deaths
  • True latent trajectories ($S,E,I,A,R$), daily $R_{t}$, and clinical outcomes
  • Intervention metadata (timing, effect size), simulation parameters, and noise specifications
This end-to-end framework yields a high-diversity corpus of simulated epidemics with realistic mechanistic and observational properties, enabling generalizable pretraining of SGNNs across infectious disease tasks.
Parameter Distribution / Range Description
Transmission and Progression
$\beta$ $\text{Uniform}(0.10,1.00)$ Effective transmission rate per contact
$\gamma$ $\text{Uniform}(0.10,0.33)$ Recovery rate from infection
$\sigma$ $\text{Uniform}(0.20,0.40)$ Rate of progression from exposed to infectious
$p_{A}$ $\text{Uniform}(0.10,0.70)$ Fraction of cases that are asymptomatic
$\alpha$ $\text{Uniform}(0.30,1.00)$ Relative transmissibility of asymptomatic individuals
$k$ (dispersion) $\text{Uniform}(0.10,1.00)$ (in some sims) Overdispersion for super-spreading (used in $\eta\sim\text{Gamma}(k,k)$)
$s(t)$ 1–4 harmonics, amp $\sim\text{U}(0.05,0.20)$ Seasonal forcing using sinusoidal harmonics
Demography and Immunity
NN $\text{LogUniform}(5\times 10^{4},5\times 10^{7})$ Total population size
$\mu$ $\text{Uniform}(0,1/365)$ Per capita birth and death rate (if enabled)
$\omega$ $\text{Uniform}(0.001,0.0075)$ Waning immunity rate (R $\rightarrow$ S)
Importation rate $\text{LogUniform}(0.0001,0.01)$ Daily chance of importation into $E$ or II
Clinical Outcomes
Hosp. prob. $\text{Uniform}(0.02,0.15)$ Probability of hospitalization (per wave)
Death — Hosp. $\text{Uniform}(0.05,0.30)$ Probability of death conditional on hospitalization
Hosp. delay $\text{Gamma}(\mu,\theta)$; $\mu\sim\text{Uniform}(5,12)$ Delay from infection to hospitalization
Death delay $\text{Gamma}(\mu,\theta)$; $\mu\sim\text{Uniform}(14,21)$ Delay from infection to death
Intervention (NPI) Parameters
Trigger threshold $\text{Uniform}(0.001N,0.01N)$ Cases required to trigger intervention
Relaxation threshold $\text{Uniform}(0.0002N,0.002N)$ Cases required to relax intervention
Reduction factor $\text{Uniform}(0.20,0.60)$ Reduction in $\beta(t)$ during intervention
Duration $\text{Uniform}(14,120)$ (days) Minimum intervention duration
Observation Model (Reporting)
Initial report rate $\text{Uniform}(0.05,0.40)$ Case reporting probability early in epidemic
Final report rate $\text{Uniform}(0.25,0.85)$ Case reporting probability later in epidemic
Logistic midpoint $\text{Uniform}(0.2,0.7)$ of time axis Center point for logistic increase in reporting
Delay kernel mode $\text{Uniform}(0,3)$ (days) Mode of geometric distribution for reporting delay
Weekday effects $\text{Normal}(1.0,0.05)$ Multiplicative weekday effects (e.g., weekend drop)
Case noise $\text{LogNormal}(0,\sigma^{2}),\sigma\sim\text{U}(0.15,0.25)$ Multiplicative noise applied to case counts
Hosp noise $\text{LogNormal}(0,\sigma^{2}),\sigma\sim\text{U}(0.10,0.15)$ Noise applied to hospitalization counts
Death noise $\text{LogNormal}(0,\sigma^{2}),\sigma\sim\text{U}(0.05,0.10)$ Noise applied to death counts
Table 1: Summary of stochastic simulation parameter distributions used to generate mechanistically diverse, observationally realistic synthetic outbreaks.
A.2 Ecological Simulation Frameworks
A.2.1 Butterfly Simulation Framework
Theoretical model.
To generate synthetic multivariate population dynamics for the butterfly forecasting experiments, we developed a hybrid mechanistic simulation framework that blends ecological growth models, interspecific competition, environmental stochasticity, and realistic observation processes. This framework is designed to approximate both the underlying biological drivers of species abundance fluctuations and the statistical properties of survey-based monitoring data.
Each simulation begins by sampling a random number of species (up to 32), with species-specific growth parameters drawn independently. The intrinsic growth rate for species $i$, denoted $r_{i}$, is drawn from a uniform distribution $r_{i}\sim\mathcal{U}(0.15,0.4)$. Baseline carrying capacities $K_{i}$ are assigned as scaled multiples of initial population sizes, with $K_{i}=N_{0i}\times U(1.5,2.5)$, where $N_{0i}\sim 10^{\mathcal{U}(1.7,2.4)}$. This yields baseline abundances ranging from approximately 50 to 250 individuals per species. Interspecific competition is modeled using a Lotka-Volterra framework [36], with pairwise competition coefficients $\alpha_{ij}$ drawn from a truncated normal distribution $\alpha_{ij}\sim\mathcal{N}(0.03,0.01)$, clipped to be non-negative. Self-competition terms $\alpha_{ii}$ are set to zero. This results in weak but widespread competition, consistent with the diffuse resource-mediated interactions observed in real butterfly communities.
Environmental variability is introduced through two distinct mechanisms. First, longitudinal environmental shocks are modeled via an annual AR(1) process that generates temporally correlated environmental multipliers. The initial value is drawn from $E_{0}\sim\mathcal{N}(1.0,0.05)$, and subsequent years evolve as $E_{t}=0.7\cdot E_{t-1}+\epsilon_{t}$, where $\epsilon_{t}\sim\mathcal{N}(0,0.05)$. This process is exponentiated and normalized to yield multiplicative factors applied to carrying capacities over time. Second, seasonal forcing is applied via sinusoidal modulation of carrying capacities to simulate annual cycles. Specifically, we define seasonally adjusted carrying capacity as:
$K_{it}=K_{i}\cdot E_{t}\cdot\left[1+A_{s}\sin\left(2\pi\frac{t+\phi}{12}\right)\right],$
where the seasonal amplitude $A_{s}=0.15$ and the phase offset $\phi$ is drawn from $\mathcal{U}(0,1)$.
Real-world effects.
After integrating the deterministic dynamics, process noise is applied to simulate intrinsic demographic stochasticity. For each species and time point, we perturb the latent population size with Gaussian noise centered at zero and with standard deviation equal to 3% of the population value. These noisy dynamics are generated by solving a system of ordinary differential equations that govern coupled species growth and competition. For species $i$, the governing equation is:
$\frac{dN_{i}}{dt}=r_{i}N_{i}\left(1-\frac{N_{i}+\sum_{j}\alpha_{ij}N_{j}}{K_{it}}\right),$
where $K_{it}$ reflects the time-varying carrying capacity influenced by environmental and seasonal variation. The system is integrated numerically over a 100-year simulation horizon using a time step of one year via ‘scipy.integrate.odeint‘.
To simulate realistic observational data, we apply a multi-stage noise process to the latent population trajectories. First, each true abundance $N_{it}$ is converted to an observed count using a negative binomial distribution with high dispersion: $y_{it}\sim\mathrm{NB}(\mu=N_{it},\text{overdispersion}=2000)$. This models sampling error and overdispersion typical in ecological field surveys. The resulting counts are then transformed to log base 10 and perturbed with additive Gaussian noise of standard deviation 0.08 to simulate additional measurement error, rounding artifacts, and variability introduced during reporting and digitization.
The final output consists of multivariate, log-transformed species abundance time series, representing realistic synthetic training data for ecological forecasting tasks. These data reflect complex population dynamics shaped by nonlinear growth, weak competition, temporally correlated environmental shocks, seasonal cycles, demographic noise, sampling error, and layered observational uncertainty.
A.2.2 Lynx-Hare Simulation Framework
To simulate predator-prey cycles reflective of historical lynx-hare dynamics, we implemented a stochastic extension of the Rosenzweig-MacArthur model [21], incorporating a mild self-limitation term on predator growth. The model consists of two coupled differential equations describing the dynamics of the hare ($H$) and lynx ($L$) populations:
$\frac{dH}{dt}=rH\left(1-\frac{H}{K}\right)-\beta HL$
$\frac{dL}{dt}=\delta HL-\gamma L-\rho L^{2},$
where $r$ is the intrinsic growth rate of hares, $K$ is the hare carrying capacity, $\beta$ is the predation rate, $\delta$ is the conversion efficiency from prey to predator, $\gamma$ is the natural mortality rate of lynx, and $\rho$ is a density-dependent predator self-limitation term. To ensure stable oscillatory dynamics, parameters are sampled from broad but biologically plausible ranges: $r\sim\mathcal{U}(0.4,0.6)$, $K\sim\mathcal{U}(80,120)$, $\beta\sim\mathcal{U}(0.02,0.04)$, $\delta\sim\mathcal{U}(0.025,0.04)$, $\gamma\sim\mathcal{U}(1.0,2.0)$, and $\rho\sim\mathcal{U}(0.0005,0.002)$.
Stochasticity is introduced during forward iteration by injecting Gaussian demographic noise into both populations at each time step. The noise magnitude is scaled inversely with population size, reflecting the empirical observation that larger populations exhibit lower relative variance. Specifically, the coefficients of variation are computed as:
$\mathrm{CV}_{H}=\min\left(0.1,\frac{200}{1000\cdot H}\right),\quad\mathrm{CV}_{L}=\min\left(0.1,\frac{100}{1000\cdot L}\right).$
Population updates are drawn from normal distributions with these variances, while ensuring all values remain non-negative and bounded by ecologically plausible maxima ($H_{\max}=200$, $L_{\max}=80$).
Final simulation outputs consist of annual time series for hare and lynx populations over a 100-year period. To match the structure of real-world empirical records such as fur trapping data, outputs are expressed in ”pelt-unit” scaling and passed through the same noise pipeline as the butterfly model to ensure realistic training data for low-dimensional predator-prey forecasting tasks.
A.3 Synthetic Data Generation for Suzuki-Miyaura Cross-Coupling
To support pretraining in the domain of small-molecule synthesis, we constructed a hybrid simulation framework for the Suzuki–Miyaura cross-coupling reaction [22]—a palladium-catalyzed process widely used in pharmaceutical and materials chemistry to form carbon–carbon bonds between aryl halides and boronates [22]. Our simulator was designed to produce synthetic reaction yields that match the structural patterns and failure behavior observed in the empirical dataset of Perera et al. [23], which includes 5,760 nanomole-scale reactions executed on an automated flow synthesis platform.
While our simulator does not explicitly solve chemical reaction networks via ODEs or stochastic kinetics, it is nonetheless grounded in mechanistic principles. This design reflects the nature of the available data: the experimental dataset reports only final reaction outcomes (i.e., isolated yields), not time-resolved concentration trajectories. As such, our focus is on modeling endpoint behavior rather than simulating reaction dynamics. We embed domain knowledge through structured component effects, chemically motivated interaction terms, and expert-informed failure rules that reflect known catalytic behaviors (e.g., ligand–substrate incompatibilities, base–boronate failures). The result is a scalable generative framework that captures core mechanistic dependencies without requiring first-principles simulation, enabling effective pretraining of SGNNs on high-fidelity synthetic data.
Simulation Design
We modeled reaction yield as a structured function of five categorical inputs: aryl halide, boronate, ligand, base, and solvent. For each component $x$ in category $C$, we estimated a main effect $\mu_{x}$ as:
$\mu_{x}=\mathbb{E}[\text{Yield}\mid x]-\mathbb{E}[\text{Yield}]$ (2)
These were combined to form a base prediction:
$\hat{y}_{\text{base}}=\mu+\sum_{C}\mu_{x_{C}}$ (3)
where $\mu$ is the global yield mean and $x_{C}$ is the component selected from category $C$.
To capture higher-order dependencies, we added:
  • Pairwise interactions $\delta_{x_{C},x_{C^{\prime}}}$ for combinations with sufficient empirical support (e.g., aryl halide–ligand, base–boronate).
  • Three-way interactions $\delta_{x_{A},x_{B},x_{C}}$ for known mechanistic phases (e.g., oxidative addition: aryl halide–ligand–solvent).
  • Exact memorization for all reaction tuples observed in the Perera dataset, storing their empirical mean and variance for maximal fidelity.
This yields a composite predictive model:
$\hat{y}=\hat{y}_{\text{base}}+\sum\delta_{\text{pair}}+\sum\delta_{\text{3way}}+\epsilon$ (4)
Failure Modeling
Reactions with yields below 5% were treated as failures. We estimated the probability of failure using a hybrid method:
  • Empirical failure rates computed per component.
  • Expert-driven heuristics, such as:
    • Strong base + weak ligand (e.g., NaOtBu + PPh3)
    • BF3K boronates with weak bases
    • Aryl chlorides with low-activity ligands
The final failure probability $p_{\text{fail}}$ was computed as a weighted sum of these effects. Reactions sampled as failures were assigned yields drawn from a uniform distribution in the interval $[0.001,0.04]$.
Heteroscedastic Noise and Calibration
For successful reactions, we modeled residual uncertainty using a heteroscedastic Gaussian noise model:
$)y\sim\mathcal{N}(\hat{y},\sigma^{2}(\hat{y}$)) (5)
[…]
Sampling Strategy
To ensure broad yet chemically meaningful coverage of the reaction space, we used a stratified sampling protocol:
  • 60% of reactions drawn from exact empirical combinations (memorized in the dataset)
  • 30% from high-yielding partial combinations (e.g., known aryl halide–ligand pairs paired with new bases/solvents)
  • 10% uniformly sampled across the full combinatorial space
This strategy preserves empirical structure while encouraging generalization to underexplored regions.
Dataset Summary
We generated a total of 2.5 million synthetic reactions. After calibration, the final dataset statistics were:
Metric Value
Unique synthetic reactions 2,500,000
Mean yield 0.62
Standard deviation 0.28
Failure rate (yield << 5%) 9.1%
This synthetic corpus preserves the structural complexity, sparsity, and statistical patterns of real-world reaction optimization, while enabling scalable simulation-grounded learning in low-data chemical settings.
A.4 Synthetic Cascade Generation for Source Identification
To train SGNNs for source identification, we generated synthetic diffusion cascades using the Independent Cascade (IC) model on a fixed Barabási–Albert graph [25] with 1,000 nodes, where each node adds 5 edges. The IC model captures a key property of real-world information diffusion: once activated, each node has a single chance to infect each of its neighbors with probability $p=0.05$, proceeding in discrete time steps. Each cascade begins with a randomly selected source node and unfolds for up to 15 time steps or until the process terminates.
For each simulation, we recorded the time of infection for every node. To reflect realistic observational constraints—such as incomplete contact tracing or partial social media logs—we randomly masked 20% of infected nodes by setting their infection time to a sentinel value, indicating that they were unobserved. Nodes not infected during the cascade were also assigned this sentinel value. The resulting input for each node consisted of its infection time (observed or unobserved), paired with structural encodings describing its location in the network (described in the model architecture).
This setup produces a large corpus of partially observed cascades with known ground-truth sources, enabling supervised training of SGNNs to learn latent causal structure under uncertainty.
Appendix B SGNN Architecture and Training Details
B.1 Model Architecture
To demonstrate the flexibility of the SGNN framework across tasks and domains, we implemented a range of neural architectures tailored to each problem type. We found that straightforward, task-appropriate models performed well out-of-the-box when trained with simulation-grounded supervision. This suggests that the SGNN training paradigm provides a strong inductive bias that reduces reliance on heavy architecture tuning. Accordingly, we selected architectures that align naturally with each task structure—e.g., sequence models for time series and graph models for networks—to ensure strong baseline performance without domain-specific feature engineering.
Time Series Forecasting. For epidemiological and ecological forecasting, we used two classes of models:
  • PatchTST-style temporal patching transformers [37]: These models divide input sequences into non-overlapping temporal patches (e.g., 7-day blocks) and encode each as a token, allowing efficient self-attention over longer sequences. This design captures coarse-grained temporal structure while preserving positional resolution.
  • Hybrid convolutional neural network (CNN)-Transformer architectures: We also employed models that combine convolutional feature extraction with transformer-based sequence modeling. CNN layers detect local patterns (e.g., sudden surges, weekly reporting artifacts), while transformers model long-range dependencies and dynamics. This hybrid design balances expressivity and efficiency, particularly for noisy, high-variance real-world time series with both short-term and long-term patterns.
Both architectures are capable of producing calibrated probabilistic forecasts when trained using quantile loss. Importantly, neither model uses domain-specific feature engineering or hand-crafted mechanisms—their strong performance is driven entirely by the structure learned from simulation-based pretraining.
Network Diffusion Source Identification. For source inference on partially observed diffusion cascades, we used a 4-layer Graph Attention Network (GAT) with residual connections, batch normalization, ReLU activations, and dropout. GATs are well-suited for this task because they learn to assign attention weights over neighboring nodes, enabling the model to focus on the most informative connections in a noisy or incomplete cascade.
Each node’s input consists of its observed infection time (or a sentinel value if unobserved) concatenated with its Laplacian positional encoding (LapPE), derived from the first 16 nontrivial eigenvectors of the graph Laplacian. These encodings provide global structural context and allow the model to distinguish between nodes with identical local neighborhoods but different roles in the global network topology—critical for inferring cascade origins. The GAT was trained to output a per-node probability distribution over possible sources, using cross-entropy loss.
Chemical Yield Prediction. For regression on simulated chemical reaction yields, we used a hybrid wide-and-deep SGNN architecture tailored to discrete combinatorial inputs. The model is designed to capture both linear interactions and higher-order nonlinear effects among reaction components.
  • Inputs: Each reaction is represented by five categorical variables: aryl halide, boronate, ligand, base, and solvent. Each component is embedded in a 128-dimensional learnable vector space, and the resulting embeddings are concatenated into a single 640-dimensional feature vector.
  • Wide pathway: A linear layer maps the concatenated embeddings directly to a scalar prediction. This “wide” component captures additive main effects and is designed to quickly learn interpretable marginal trends across the combinatorial design space.
  • Deep pathway: A parallel MLP first passes the concatenated embeddings through a non-linear interaction layer, followed by a series of hidden layers (768 → 768 → 384 → 128 → 64) with ReLU activations, batch normalization, and dropout. This deep component captures complex, nonlinear interactions between reaction components—such as synergy or incompatibility between specific reagents.
  • Multi-head output: The deep representation is fed into three heads:
    1. 1.
      A main regression head (yield prediction), which is summed with the wide pathway output.
    2. 2.
      An auxiliary head, providing an independent yield prediction to encourage representation diversity.
    3. 3.
      A confidence head, which outputs a [0,1] scalar using a sigmoid activation to represent model uncertainty.
  • Ensembling: The final prediction is a learnable weighted ensemble of the main and auxiliary heads. Ensemble weights are learned during training via a softmax-normalized two-element parameter vector.
This architecture is trained entirely on synthetic reaction data generated from simulation-based rulesets. All models use mean squared error loss on yield values and are evaluated in a zero-shot fashion on real reactions not seen during pretraining. The SGNN framework allows the model to learn a latent representation of reactivity space grounded in mechanistic constraints, without requiring experimental supervision.
B.2 Training Hyperparameters
All models were trained using the AdamW optimizer [38] with a weight decay of 0.01. Learning rates followed a cosine annealing schedule throughout training. Batch size was 256 for ecology, chemical yield prediction, and source identification, and 3,072 for disease forecasting (since the network was larger). The number of pretraining steps ranged from several hundred (ecology) to 10,000 steps (infectious disease forecasting). No fine-tuning was performed for any downstream evaluations; all results presented reflect direct zero-shot application of the pretrained SGNN models without additional task-specific retraining, except for in chemical yield prediction. Here, to demonstrate how fine-tuning could be applied, we pretrained on the simulated dataset and then fine-tuned on the real, smaller dataset [23]. Fine-tuning led to a 5% increase in performance (fine-tuned performance shown in figure 3b).
B.3 Loss Functions
For forecasting tasks, models were trained using quantile loss to produce probabilistic forecasts across multiple quantile levels. For scientific parameter inference tasks (e.g., estimating $R_{0}$), quantile loss was also used to generate full predictive intervals rather than point estimates. For classification tasks (source identification), cross-entropy loss was used. No auxiliary losses were applied in any experiments.
Appendix C Datasets and Evaluation
We evaluated SGNNs across diverse forecasting, inference, and regression tasks using real-world datasets from epidemiology, chemical kinetics, and ecology.
C.1 Infectious Disease Datasets and Evaluation
COVID-19 Forecasting.
We evaluated SGNNs on weekly COVID-19 mortality forecasting using data from the CDC COVID Forecast Hub [15], covering all 50 U.S. states, Washington D.C., and Puerto Rico. The evaluation period spanned April 2020 through November 2021, with weekly aggregation performed using CDC epidemiological weeks (epiweeks). SGNN models were applied zero-shot, without retraining or access to real COVID-19 data during training. Forecasting was conducted at horizons of 1 to 4 weeks ahead. Baseline models were evaluated over the same period, and all metrics were computed using the full real-world time series for each location. No normalization or transformations were applied.
Dengue Forecasting.
Dengue forecasting was conducted using weekly case data from Brazilian states, sourced from Project Tycho [18]. The SGNN was evaluated zero-shot, having been pretrained only on non-vector-borne simulations. Baseline models were trained directly on the dengue time series. For these baselines, we used a rolling-window evaluation strategy: models were retrained at each forecasting time point using all available data up to that point, to provide the strongest possible comparison. Forecast performance was then evaluated at 1–4 week horizons beyond each training window, and results were aggregated across all forecast periods and states.
$R_{0}$ Estimation.
For synthetic evaluation of reproduction number ($R_{0}$) inference, we generated a held-out test set of 5,000 simulated epidemic trajectories using mechanistic models distinct from the pretraining corpus. Ground truth $R_{0}$ values were analytically computed via next-generation matrix methods, and estimation accuracy was evaluated using mean squared error and mean percentage error. For real-world validation, we evaluated SGNN and baseline methods on early COVID-19 case trajectories from March 2020 [15]. Specifically, we used three weeks of daily reported cases per location (U.S. states, Washington D.C., and New York City) and compared SGNN $R_{0}$ estimates to values from the literature [34, 35]. These real-world $R_{0}$ values were sourced from retrospective analyses using later data, providing a plausibility check against which early-estimate performance could be judged.
Back-to-Simulation Attribution.
To evaluate SGNN interpretability, we used back-to-simulation attribution on real COVID-19 mortality time series in Michigan. The input sequence was passed through the SGNN encoder to extract its embedding. Cosine similarity was then computed between this embedding and the embeddings of 10,000 synthetic simulations used during pretraining. The top-50 most similar simulations were retrieved, and the distributions of their underlying mechanistic parameters were analyzed. This allowed us to examine whether SGNN’s internal representation aligned real-world data with mechanistically plausible synthetic regimes.
C.2 Ecological Datasets and Evaluation
Lynx-Hare Forecasting.
We evaluated SGNN on the canonical predator-prey dynamics of the Canadian lynx and snowshoe hare, using annual pelt-trapping records from 1821 to 1934 [19]. The first 40 years of data were used for fitting mechanistic and statistical baselines and served as context input for SGNN. Models forecasted population counts 4 years ahead at each evaluation point. Evaluation metrics were averaged across forecast windows beyond the initial 40-year context.
Butterfly Population Forecasting.
We evaluated SGNNs on high-dimensional ecological forecasting using the UK Butterfly Monitoring Scheme dataset [20], which provides long-term annual abundance indices for dozens of butterfly species. Data through 2023 were obtained from the UKBMS public portal. Forecasting tasks were defined at multiple dimensionalities: for each value (e.g., 5, 10, 15 species), we selected species alphabetically (e.g., first 5, then next 5, etc.). Baseline models were trained directly on the real data, using the first 80% of each species’ time series for training and the final 20% for testing. SGNN was pretrained on synthetic simulations of multi-species population dynamics and applied zero-shot. No species from the real-world dataset were included in synthetic pretraining.
C.3 Chemical Kinetics Dataset and Evaluation
We evaluated SGNNs on reaction yield prediction for palladium-catalyzed Suzuki-Miyaura cross-coupling, using the nanomole-scale dataset developed by Perera et al. [23] and accessed via the Open Reaction Database. The dataset consists of 5,760 real-world reactions screened under varying combinations of aryl halides, boronates, ligands, bases, and solvents, with yields reported as continuous values in the interval $[0,1]$.
Each reaction was represented by five categorical input features corresponding to its component identities. No continuous descriptors (e.g., temperature or concentration) were used, and reactions were modeled purely based on discrete structural combinations. Reactions with missing yield values were excluded. The dataset was randomly split into 75% for analysis (training + validation) and 25% for held-out evaluation. All reported performance results use this final test set, which was never seen during training or pretraining.
After pretraining, the SGNN was evaluated on the held-out 20% of real reactions. Model performance was assessed using the coefficient of determination ($R^{2}$), following the convention in prior work [24]. The SGNN achieved a final $R^{2}$ of 0.90, outperforming previous state-of-the-art models trained directly on real data (which achieved $R^{2}=0.85$), corresponding to a one-third reduction in residual (unexplained) variance.
Appendix D Back-to-Simulation Attribution
To enable mechanistically interpretable diagnostics, SGNNs support back-to-simulation attribution, a retrieval-based method that identifies the most similar synthetic regimes to a given real-world input based on the SGNN’s internal representation space. This section describes the implementation in full detail.
Embedding Extraction.
To perform attribution, a real-world time series (e.g., weekly COVID-19 deaths in Michigan) is first transformed into the format expected by the pretrained SGNN encoder. This involves applying the exact same preprocessing used during model training—for example, in our case, log1p transformation followed by z-score normalization using training-set statistics. The model-specific context length is enforced, with earlier time points padded if necessary. The processed input is then passed through the full SGNN encoder, which, in this case, includes multi-scale convolutional feature extraction, categorical embeddings (e.g., disease type, time-series type, calendar indices), and a CNN-Transformer backbone. The encoder output is pooled using a validity mask over the time dimension to yield a fixed-length embedding $z\in\mathbb{R}^{d}$ (with $d=1024$ in our primary implementation). While preprocessing steps and model architecture can vary by domain, attribution requires that embeddings be computed using the same pipeline and parameters as used during training.
Simulation Embedding Corpus.
To enable attribution, we construct an embedding database by encoding a large set of synthetic trajectories from the SGNN pretraining corpus. Each simulation is processed using the same data pipeline as real-world inputs, including identical preprocessing steps (e.g., normalization, padding, feature construction) and the same trained SGNN encoder. The result is a set of embedding vectors $\{z_{i}\}$, each corresponding to a simulated trajectory. Alongside each embedding, we store the full set of underlying mechanistic parameters used to generate that simulation (e.g., transmission rates, population size, environmental forcing strength), which serve as the basis for interpretation. To support efficient reuse, all embeddings and parameter metadata are precomputed and cached to disk. While the specific preprocessing steps vary across domains, consistency between simulation and real-world input processing is essential for meaningful attribution.
Similarity Search.
Given a real-world input embedding $z$, we compute cosine similarity against all stored simulation embeddings $\{z_{i}\}$:
$\text{sim}(z,z_{i})=\frac{z\cdot z_{i}}{\|z\|\|z_{i}\|}.$
We retrieve the top-$k$ simulations with highest similarity scores (typically $k=50$) using exact brute-force similarity via NumPy or scikit-learn. In practice, cosine similarity computation across 10,000+ embeddings completes in under 5 seconds on a single CPU core.
Mechanistic Interpretation.
Each retrieved simulation is associated with its full parameter dictionary $\theta_{i}$. We analyze the top-$k$ simulations to inspect the distributions of their mechanistic parameters. For example, when applied to Michigan COVID-19 death data, the top retrieved simulations cluster tightly around known population size ($\sim$10 million), plausible hospitalization and death probabilities, and realistic asymptomatic proportions. This validates that the model’s internal embedding space captures meaningful system-level dynamics and aligns real inputs with plausible mechanistic regimes.
Caching and Reproducibility.
To accelerate repeated attribution queries, we cache both the simulation embeddings and parameter metadata as serialized binary files. These are loaded once per session and used for all future retrievals. If the cache does not exist, simulation embeddings are computed on demand and saved to disk.
Scalability.
The full back-to-simulation attribution pipeline is implemented in PyTorch and NumPy and runs efficiently on CPU/GPU. Embedding a real input series requires only a forward pass through the encoder (under 100 ms on GPU), while similarity search over 10,000+ simulation embeddings takes under 5 seconds on CPU. This enables interactive diagnostic use cases and large-scale attribution across real-world datasets.
Appendix E Additional Results
E.1 Ecological Forecasts
To illustrate the forecasting capabilities of SGNNs in ecological systems, we provide a visual comparison of model predictions against ground-truth population dynamics in the lynx-hare predator-prey system. Figure 5 shows forecast trajectories from SGNNs, Rosenzweig-MacArthur growth model (RMG), a purely mechanstic model, and vector autoregressive moving average models (VARMA), a purely data-driven model, across four rolling prediction windows.
These results emphasize the capability of simulation-grounded pretraining for capturing complex ecological feedbacks and suggest that SGNNs are capable of robust extrapolation even in systems with limited historical data and strong cyclical behavior.
Figure 5: SGNNs outperform classical models in forecasting lynx-hare predator-prey dynamics. Rolling forecasts are shown for four evaluation windows, comparing SGNN (red) to mechanistic models (RMG, blue) and VARMA (green) models. True population trajectories are plotted in black. SGNNs consistently maintain accurate phase and amplitude tracking, while baselines deteriorate.
E.2 State $R_{0}$ Estimates
To further evaluate SGNNs ability to infer unobservable epidemiological parameters, we estimated the basic reproduction number ($R_{0}$) for early COVID-19 outbreaks across all U.S. states and territories. Estimates were generated using SGNNs trained solely on synthetic data, and applied in a fully zero-shot setting to early outbreak curves (February–March 2020), consisting of three weeks of reported case data per location. Confidence intervals were taken from the quantile predictions of the SGNN.
SGNN-estimated early COVID-19 $R_{0}$ values across U.S. states and territories. Estimates reflect SGNN inference from early outbreak time series, using a model trained exclusively on synthetic simulation data. Confidence intervals denote 95% quantile-based uncertainty.
Location $R_{0}$ (95% CI)
AK 3.77 (3.19–6.06)
AL 3.24 (2.08–3.80)
AR 1.97 (1.53–2.52)
AS 1.13 (0.90–1.27)
AZ 2.09 (1.90–2.91)
CA 4.95 (4.58–5.54)
CO 1.46 (1.06–1.74)
CT 1.10 (1.03–1.21)
DC 1.87 (1.33–1.98)
DE 1.59 (0.95–1.75)
FL 1.68 (1.22–2.06)
FSM 0.96 (0.80–1.27)
GA 3.06 (2.72–3.13)
GU 3.26 (2.96–5.63)
HI 4.11 (3.03–5.37)
IA 2.06 (1.45–2.90)
ID 3.76 (2.43–5.80)
IL 3.41 (2.86–3.55)
IN 2.74 (1.97–3.00)
KS 1.53 (1.17–1.71)
KY 1.94 (1.45–2.49)
LA 4.12 (3.86–5.31)
MA 4.68 (4.05–5.69)
MD 1.16 (0.82–1.73)
ME 1.30 (1.28–1.73)
MI 3.49 (3.09–3.67)
MN 1.33 (1.24–2.18)
MO 2.41 (1.86–3.11)
MP 2.52 (1.73–4.21)
MS 2.13 (1.64–2.47)
Location $R_{0}$ (95% CI)
MT 4.33 (3.98–6.08)
NC 1.84 (1.37–1.92)
ND 3.47 (2.74–5.26)
NE 0.98 (0.85–1.49)
NH 1.40 (1.38–2.32)
NJ 3.02 (2.62–3.98)
NM 0.75 (0.71–1.04)
NV 4.08 (3.44–7.41)
NY 6.11 (5.10–6.74)
NYC 6.14 (5.02–7.97)
OH 2.35 (2.11–2.17)
OK 2.30 (1.64–2.77)
OR 4.63 (4.25–9.04)
PA 3.11 (2.80–3.13)
PR 3.16 (2.07–4.52)
PW 0.96 (0.80–1.27)
RI 1.23 (1.07–1.57)
RMI 0.96 (0.80–1.27)
SC 1.80 (1.27–2.09)
SD 0.89 (0.73–1.46)
TN 2.87 (2.59–3.01)
TX 2.98 (2.46–3.15)
UT 1.80 (1.38–1.95)
VA 0.97 (0.81–1.14)
VI 6.79 (4.97–9.49)
VT 2.91 (1.98–4.03)
WA 4.80 (4.25–5.00)
WI 1.59 (1.45–1.91)
WV 2.20 (1.62–2.79)
WY 2.66 (2.34–4.59)
The SGNN produced high and variable early-$R_{0}$ estimates across locations, with several important observations:
[…]
  • Low-$R_{0}$ regions and territories: FSM, PW, RMI, NM, NE, and VA all yielded $R_{0}$ values near or below 1. This likely reflects true containment in isolated areas (e.g., FSM, RMI) or data sparsity in early periods, which limits identifiability and pulls predictions toward lower priors.
  • General interpretability: SGNN’s estimates tend to reflect known drivers of early outbreak heterogeneity—population density, travel exposure, and reporting infrastructure—despite having no access to these covariates. This suggests that SGNN’s latent dynamics have internalized meaningful global patterns purely from time-series shape and signal-to-noise characteristics.
Overall, these results demonstrate SGNN’s ability to produce informative and accurate $R_{0}$ estimates directly from early outbreak curves. This capability—learning to infer latent, unobservable scientific quantities with calibrated uncertainty—would be impossible using traditional supervised learning. It highlights the potential for simulation-grounded models to support real-time scientific inference during future emergent crises.